Butane, C4H10, is a component of natural gas that is used as fuel for cigarette lighters. The balanced equation of the complete combustion of butane is

2C4H10(g)+13O2(g)→8CO2(g)+10H2O(l)

At 1.00 atm and 23 ∘C, what is the volume of carbon dioxide formed by the combustion of 1.60 g of butane?
1
Expert's answer
2019-05-06T05:07:07-0400

Solution

From the equation it can be concluded that 2 moles of butane form 8 moles of CO2. So 1 mole of butane form 4 mole of CO2. Find the amount of moles of butane given:


"n(C4H10) = m\/M = 1.6\/58 = 0.0275 (mol)"

Find the amount of CO2 formed:


"n(CO2) = 0.0275*4 = 0.11 (mol)"

If PV=nRT for each gas at any conditions, find which volume is occupied by 0.11 mol of CO2 at the conditions indicated in the task:


"V = nRT\/P = nRT"

"V = 0.11*(273+23)*0.082 = 2.67 (L)"

Answer

2.67 L of carbon dioxide are formed by the combustion of 1.60 g of butane.